# Multi-Objective Large Language Model Unlearning

### v1

@@ Abstract
Machine unlearning in the domain of large language models (LLMs) has attracted great attention recently, which aims to effectively eliminate undesirable behaviors from LLMs without full retraining from scratch. In this paper, we explore the Gradient Ascent (GA) approach in LLM unlearning, which is a proactive way to decrease the prediction probability of the model on the target data in order to remove their influence. We analyze two challenges that render the process impractical: gradient explosion and catastrophic forgetting. To address these issues, we propose Multi-Objective Large Language Model Unlearning (MOLLM) algorithm. We first formulate LLM unlearning as a multi-objective optimization problem, in which the cross-entropy loss is modified to the unlearning version to overcome the gradient explosion issue. A common descent update direction is then calculated, which enables the model to forget the target data while preserving the utility of the LLM. Our empirical results verify that MoLLM outperforms the SOTA GA-based LLM unlearning methods in terms of unlearning effect and model utility preservation.
@@ I Introduction
  • We formulate LLM unlearning as a multi-objective optimization problem and calculate a common descent direction to solve it, allowing the model to unlearn the target data while preserving the utility.
@@ III-A Problem Formulation
To balance the trade-off between unlearning and model utility preservation, we formulate LLM unlearning as the following multi-objective optimization problem:
@@
III-B Common Descent Update Direction
@@
We propose a novel and efficient way to determine such a common descent direction $d^{t}$. Let $S$ be the space spanned by $g_{fgt}$, $g_{KL}$, and $g_{rt}$. The key idea is to obtain the dual vectors of $g_{fgt}$, $g_{KL}$, $g_{rt}$, denoted as $g_{fgt}^{*}$, $g_{KL}^{*}$, $g_{rt}^{*}$, where the convex combinations of $g_{fgt}^{*}$, $g_{KL}^{*}$, $g_{rt}^{*}$ form the dual space of $S$, as seen in Fig. 1. Afterward, the common descent direction $d^{t}$ for the model update can be obtained by taking the average of $-g_{fgt}^{*}$, $-g_{KL}^{*}$, and $-g_{rt}^{*}$.

### v2

@@ Abstract
Machine unlearning in the domain of large language models (LLMs) has attracted great attention recently, which aims to effectively eliminate undesirable behaviors from LLMs without full retraining from scratch. In this paper, we explore the Gradient Ascent (GA) approach in LLM unlearning, which is a proactive way to decrease the prediction probability of the model on the target data in order to remove their influence. We analyze two challenges that render the process impractical: gradient explosion and catastrophic forgetting. To address these issues, we propose Multi-Objective Large Language Model Unlearning (MOLLM) algorithm. We first formulate LLM unlearning as a multi-objective optimization problem, in which the cross-entropy loss is modified to the unlearning version to overcome the gradient explosion issue. A common descent update direction is then calculated, which enables the model to forget the target data while preserving the utility of the LLM. Our empirical results verify that MoLLM outperforms the SOTA GA-based LLM unlearning methods in terms of unlearning effect and model utility preservation. The source code is available at https://github.com/zibinpan/MOLLM.
@@ I Introduction
  • We formulate LLM unlearning as a multi-objective optimization problem and propose a Dual Space Multiple Gradient Descent Algorithm (DS-MGDA) to calculate a common descent update direction, allowing the model to unlearn the target data while preserving the utility.
@@ III-A Problem Formulation
To balance the trade-off between unlearning and model utility preservation, we formulate LLM unlearning as the following multi-objective optimization problem (MOP):
@@
III-B Dual Space Multiple Gradient Descent Algorithm
@@
We propose a novel and efficient way to determine such a common descent direction $d^{t}$, named Dual Space Multiple Gradient Descent Algorithm (DS-MGDA). Let $S$ be the space spanned by $g_{fgt}$, $g_{KL}$, and $g_{rt}$. The key idea is to obtain the dual vectors of $g_{fgt}$, $g_{KL}$, $g_{rt}$, denoted as $g_{fgt}^{*}$, $g_{KL}^{*}$, $g_{rt}^{*}$, where the convex combinations of $g_{fgt}^{*}$, $g_{KL}^{*}$, $g_{rt}^{*}$ form the dual space of $S$, as seen in Fig. 1. Afterward, the common descent direction $d^{t}$ for the model update can be obtained by taking the average of $-g_{fgt}^{*}$, $-g_{KL}^{*}$, and $-g_{rt}^{*}$.
@@
DS-MGDA is efficient, as it relies solely on fundamental matrix computations. Moreover, it is applicable to MOP with more than 3 objectives, provided that the matrix $A$ is full-rank.
@@ Acknowledgment
Compliance with Ethical Standards
This research study was conducted using data made available in open access by (https://github.com/PKU-Alignment/safe-rlhf). Ethical approval was not required as confirmed by the license attached with the open access data.